# Yang-Lee Quantum Criticality in Various Dimensions

### v1

@@ 7 Future Directions
Our work provide useful input for future conformal bootstrap study. For non-hermitian CFTs it is difficult to use the conventional convex optimization setup, but setups based on solving truncated crossing equations[31, 32, 33] can still be applied. Without positivity there is no rigorous error bar, so one need to start with a heuristic guess of the qualitative structure of the solution, and the CFT data we extract will be useful in this aspect. Remarkably, there are recent works[82, 83] of 5-point function bootstrap using the truncated crossing equation method have successfully extracted spin OPE coefficients such as $C_{\phi TT}$ for 3D Ising model. Such quantities would take bootstrap setup of mixed $\phi$-$T$ 4-point functions system, which was recently done in [84]. Combining our results with these bootstrap techniques will hopefully significantly improve our understanding of non-hermitian CFTs.
@@ Appendix B Constructing Spinning Correlation Functions in Fuzzy Sphere
$V_{i,jk}$ $=\frac{\left(x_{i}-x_{j}\right)\cdot\left(x_{i}-x_{j}\right)\left(x_{i}-x_{k}\right)\cdot z_{i}-\left(x_{i}-x_{k}\right)\cdot\left(x_{i}-x_{k}\right)\left(x_{i}-x_{j}\right)\cdot z_{i}}{\left(x_{j}-x_{k}\right)\cdot\left(x_{j}-x_{k}\right)}$
@@
$|{\cal O}_{\ell m}\rangle$ $=\mathbf{v}_{\ell,\ell,m}^{\mu_{1},\mu_{2}\cdots\mu_{\ell}}\lim_{x\rightarrow 0}{\cal O}_{\mu_{1}\mu_{2}\cdots\mu_{\ell}}(x)|0\rangle$ (B.5)
@@
where the tensor $\mathbf{v}_{n,\ell,m}$ projects the tensor product of $n$ indices to the $(\ell,m)$ representation. The vector is normalized to $\mathbf{v}\cdot\mathbf{v}=1$ A construction of such vectors is shown below
@@
where $\mathcal{R}:f^{\mu_{1}\mu_{2}\cdots\mu_{n}}\mapsto f^{\mu_{n}\mu_{1}\cdots\mu_{n-1}}$ generates the cyclic permutation. $\big{\langle}|\ell-1|,m-m^{\prime};1,m^{\prime}\big{|}\ell,m\big{\rangle}$ is the Clebsch-Gordan coefficient. The ket-state is obtained from spacial inversion [9]
$\langle{\cal O}_{\ell m}|$ $=\mathbf{v}_{\ell,\ell,m}^{\mu_{1},\mu_{2}\cdots\mu_{\ell}}\lim_{x\rightarrow 0}x^{-2\Delta_{\cal O}}I_{\mu_{1}\nu_{1}}(x)I_{\mu_{2}\nu_{2}}(x)\cdots I_{\mu_{\ell}\nu_{\ell}}(x){\cal O}^{\nu_{1}\nu_{2}\cdots\nu_{\ell}}\left(\frac{x}{|x|^{2}}\right)|0\rangle$ (B.7)
@@
$\langle{\cal O},\ell|n_{s}^{a}|{\cal O}^{\prime},\ell^{\prime}\rangle$ $=\delta_{{\cal O},{\cal O}^{\prime}}\delta_{s,0}b_{0}+b_{a,1}f_{\ell,\ell^{\prime},s}C_{{\cal O}\phi{\cal O}^{\prime}}^{(s)}+\cdots$ (B.11)
$f_{\ell,\ell^{\prime},s}$ $=\alpha_{s}\frac{\big{(}\mathbf{v}_{s,s,0}\cdot D_{x_{2}}^{\ell}\big{)}}{s!(1/2)_{s}}\frac{\big{(}\mathbf{v}_{\ell,\ell,0}\cdot D_{z_{1}}^{\ell}\big{)}}{\ell!(1/2)_{\ell}}\frac{\big{(}\mathbf{v}_{\ell^{\prime},\ell^{\prime},0}\cdot D_{z_{3}}^{\ell^{\prime}}\big{)}}{\ell^{\prime}!(1/2)_{\ell^{\prime}}}$
$~{}~{}~{}~{}(z_{1}\cdot z_{3})^{\frac{1}{2}(\ell_{1}+\ell_{3}-\ell_{2})}(z_{1}\cdot x_{2})^{\frac{1}{2}(\ell_{1}+\ell_{2}-\ell_{3})}(z_{3}\cdot x_{2})^{\frac{1}{2}(\ell_{2}+\ell_{3}-\ell_{1})}~{}.$

### v2

@@ 7 Future Directions
Our work also provides useful input for future conformal bootstrap studies. For non-unitary CFTs it is impossible to use the conventional convex optimization methods, but setups based on solving truncated crossing equations[42, 43, 44] can still be applied. Without positivity this method is not rigorous, so one need to start with a guess of the qualitative structure of the solution, and the CFT data we provided in this paper will be useful in this respect. Remarkably, there are recent works [94, 95] on 5-point function bootstrap using the truncated crossing equation method, which successfully extracted spin OPE coefficients such as $C_{\phi TT}$ for the 3D Ising model. Such quantities can also be obtained from bootstrap of the mixed $\phi$-$T$ 4-point functions system, which was recently done in [96]. Combining our results with these bootstrap techniques can hopefully improve our understanding of non-unitary CFTs.
@@ Appendix B Constructing Spinning Correlation Functions in Fuzzy Sphere
$V_{i,j,k}$ $=\frac{\left(x_{i}-x_{j}\right)\cdot\left(x_{i}-x_{j}\right)\left(x_{i}-x_{k}\right)\cdot z_{i}-\left(x_{i}-x_{k}\right)\cdot\left(x_{i}-x_{k}\right)\left(x_{i}-x_{j}\right)\cdot z_{i}}{\left(x_{j}-x_{k}\right)\cdot\left(x_{j}-x_{k}\right)}$
@@
$|{\cal O}_{\ell m}\rangle$ $=v_{\ell,\ell,m}^{\mu_{1}\mu_{2}\cdots\mu_{\ell}}\lim_{x\rightarrow 0}{\cal O}_{\mu_{1}\mu_{2}\cdots\mu_{\ell}}(x)|0\rangle$ (B.5)
@@
where the tensor $\mathbf{v}_{n,\ell,m}$ ($v_{\ell,\ell,m}^{\mu_{1}\mu_{2}\cdots\mu_{\ell}}$ being the component form) projects the tensor product of $n$ indices to the $(\ell,m)$ representation. The vector is normalized to $\mathbf{v}\cdot\mathbf{v}=1$. A construction of such vectors is shown below
@@
where $\mathcal{R}:f^{\mu_{1}\mu_{2}\cdots\mu_{n}}\mapsto f^{\mu_{n}\mu_{1}\cdots\mu_{n-1}}$ generates the cyclic permutation and $\mathcal{R}^{k}$ is permuting $k$ times. $\big{\langle}|\ell-1|,m-m^{\prime};1,m^{\prime}\big{|}\ell,m\big{\rangle}$ is the Clebsch-Gordan coefficient. The ket-state is obtained from spacial inversion [9]
$\langle{\cal O}_{\ell m}|$ $=v_{\ell,\ell,m}^{\mu_{1},\mu_{2}\cdots\mu_{\ell}}\lim_{x\rightarrow 0}x^{-2\Delta_{\cal O}}I_{\mu_{1}\nu_{1}}(x)I_{\mu_{2}\nu_{2}}(x)\cdots I_{\mu_{\ell}\nu_{\ell}}(x){\cal O}^{\nu_{1}\nu_{2}\cdots\nu_{\ell}}\left(\frac{x}{|x|^{2}}\right)|0\rangle$ (B.7)
@@
$\langle{\cal O},\ell|M_{s}^{a}|{\cal O}^{\prime},\ell^{\prime}\rangle$ $=\delta_{{\cal O},{\cal O}^{\prime}}\delta_{s,0}b_{0}+b_{a,1}f_{\ell,\ell^{\prime},s}C_{{\cal O}\phi{\cal O}^{\prime}}^{(s)}+\cdots$ (B.11)
$f_{\ell,\ell^{\prime},s}$ $=\alpha_{s}\frac{\big{(}\mathbf{v}_{s,s,0}\cdot D_{x_{2}}^{s}\big{)}}{s!(1/2)_{s}}\frac{\big{(}\mathbf{v}_{\ell,\ell,0}\cdot D_{z_{1}}^{\ell}\big{)}}{\ell!(1/2)_{\ell}}\frac{\big{(}\mathbf{v}_{\ell^{\prime},\ell^{\prime},0}\cdot D_{z_{3}}^{\ell^{\prime}}\big{)}}{\ell^{\prime}!(1/2)_{\ell^{\prime}}}$
$~{}~{}~{}~{}(z_{1}\cdot z_{3})^{\frac{1}{2}(\ell_{1}+\ell_{3}-s)}(z_{1}\cdot x_{2})^{\frac{1}{2}(\ell_{1}+s-\ell_{3})}(z_{3}\cdot x_{2})^{\frac{1}{2}(s+\ell_{3}-\ell_{1})}~{}.$